OpenStudy (anonymous):

Find the value of g(f(2) + 4), where f and g are the following functions. g(t)=t^2-t and f(x)=1+x

OpenStudy (amistre64):

this latest problem? :)

OpenStudy (anonymous):

yes sir

OpenStudy (amistre64):

looks like we should find an f(2) first off, what do you get for f(2)?

OpenStudy (anonymous):

So plug in 2 for both?

OpenStudy (amistre64):

just for f(x) ... nothing else at the moment

OpenStudy (anonymous):

so 3

OpenStudy (amistre64):

good; now we can use that: g(f(2)+4) = g(3+4) = g(7) what do you get for g(7) ?

OpenStudy (anonymous):

42

OpenStudy (amistre64):

yep, thats right